OpenStudy (anonymous):

HELP PLEASE During a space shuttle launch, a maneuver is scheduled to begin at T minus 80 seconds, which is 80 seconds before liftoff. The maneuver lasts 1 1/2 minutes. At what time will the maneuver be complete?

ganeshie8 (ganeshie8):

It starts at time : \(T-80\) seconds and lasts for `1 1/2` minutes

ganeshie8 (ganeshie8):

how many seconds are there in `1 1/2` minutes ?

OpenStudy (anonymous):

90

ganeshie8 (ganeshie8):

add that to the `start time` to get the completing time

ganeshie8 (ganeshie8):

completing time = \(T - 80 + 90\) = ?

OpenStudy (anonymous):

so add that to 80?

OpenStudy (anonymous):

170? is that when the maneuver will be complete?

ganeshie8 (ganeshie8):

try again, whats the value of \(T-80+90\) ?

OpenStudy (anonymous):

T doesnt have a value it jus meants the amounbt of seconds before launch

ganeshie8 (ganeshie8):

Exactly ! \(T\) is the `liftoff time`, and the maneuver starts\(80\) seconds before and it takes \(90\) seconds to finish, So the completing time = \(T - 80 + 90 = T + 10\) seconds

ganeshie8 (ganeshie8):

That means the maneuver completes exactly \(10\) seconds after the `liftoff` ! if that makes any sense...

OpenStudy (anonymous):

yes!! 10 was the first answer i got when i did it by myself

OpenStudy (anonymous):

thanx

ganeshie8 (ganeshie8):

Great ! the answer is \(T + 10\) right ? (10 seconds after the liftoff)

OpenStudy (anonymous):

yes it was